One positive number is 5 times another positive number. The difference between the two numbers is 148. Find the numbers.
Answer by checkley77(12844)

x=5y
x-y=148
5y-y=148
4y=148
y=148/4
y=37 answer
x=5*37
x=185 answer
Proof:
185-37=148
148=148